OpenStudy (anonymous):

Find the arc length of the graph of the function f(x)=5 \sqrt{x^{3}} from x= 5 to x=6

OpenStudy (anonymous):

so i know the arclength formula is \[\int\limits_{5}^{6}\sqrt{1+f'(x)^2}\]

OpenStudy (anonymous):

and f(x)=5x^(3/2) f'(x)=\[15/2\sqrt{x}\]

OpenStudy (anonymous):

i get to \[\int\limits_{5}^{6}\sqrt{1+225/4(x)}dx\]

myininaya (myininaya):

do a substitution on the inside of the sqrt u=1+225x/4

OpenStudy (anonymous):

du=225/4dx?

myininaya (myininaya):

yep 4du/225=dx

myininaya (myininaya):

Don't forget to change the limits too.

OpenStudy (anonymous):

then you evaluate your a and b with u

myininaya (myininaya):

\[\frac{4}{225} \int\limits_{1+\frac{225(5)}{4}}^{1+\frac{225(6)}{4}}\sqrt{u} du\]

OpenStudy (anonymous):

does it then go to \[4/225\int\limits_{282.25}^{338.5}\sqrt{u}du\]

OpenStudy (anonymous):

to\[4/225(2u ^{3/2}/3)\left| \right|\]

OpenStudy (anonymous):

from a->b

myininaya (myininaya):

last part is to plug in and subtract

OpenStudy (anonymous):

thanks